# reference.bib

### v2

@@ 4 Landscape analysis of the objective function
In Section 4.1 and Section 4.2 respectively, we provide the landscape for these choices for two different canonical examples (Sections 2.3.1 and 2.3.2).
4.1 Two Stream example
We report results for the two-stream instability problem in this section.
For Choice A, with one parameter, we conduct a sweeping of PDE solutions by setting $b_{1}$ to be in the re-defined range of $[-0.04,0.04]$, and compute the evaluation of the three different objective functions ($L^{2}$, KL and electric energy). The result is plotted in Figure 3. Evidently, in this 1D situation, both ($L^{2}$) and (KL) give non-convex behavior in the neighborhood of the minimum, whereas (EE) demonstrates a distinctly convex profile.
Figure 3: Landscape of the Two Stream instability on the domain $[-0.04,0.04]$ for $b_{1}$ for (4.1). The objective functions are ($L^{2}$) (left), (KL) (center), and (EE)(right).
@@ 4.1 Two Stream example
  • Far field: $[-1,1]\times[-1,1]$;
  • Mid-range: $[-0.1,0.1]\times[-0.1,0.1]$
  • Near: $[-0.003,0.001]\times[-0.003,0.001]$
The wordings are selected based on the conclusions drawn from Section 3 of dispersion relation, where it was suggested the optimal $H$ should be in the range of $O(\epsilon)$. In this case, $\epsilon=0.001$. In Figure 4, we plot the values for the three different objective functions ($L^{2}$, KL and electric energy) in these three domains, and some conclusions are drawn:
  • The landscapes for $L^{2}$ and KL exhibit very similar features, both displaying pronounced non-convex behavior in the far-field and mid-range regions. Additionally, all plots are symmetric with respect to $b_{1}$, which corresponds to the first mode.
  • The landscape close to the global minimum (near-range) is convex for all three objective functions.
  • In the mid-range set, ($L^{2}$) and (KL) present a pronounced bulge around the minimum that highlights a clear non-convex behavior along the first mode direction ($b_{1}$ variable), while (EE) continues being convex.
  • In the far field, ($L^{2}$) and (KL) are oscillatory. (EE) also presents minima for large values of $b_{1}$ and $b_{2}$. In this setting, the imposed electric field $H$ is quite large, and completely dominates the dynamics. Though the electric energy $\mathcal{E}_{f}$ is small, the physics is unclear.
Figure 4: Landscape of the Two Stream instability with on the domain $[-1,1]^{2}$ (top), $[-0.1,0.1]^{2}$ (center) and, $[-0.003,0.001]^{2}$ (bottom) for $(b_{1},b_{2})$ for (4.2). The objective functions are ($L^{2}$) (left), (KL) (center), and (EE)(right).
@@ 4.2 Bump-on-Tail example
In this section, we present results for the bump-on-tail instability problem.
For Choice A, we perform a parameter sweep of the PDE solutions by varying $b_{1}$ over the redefined interval $[-0.07,0.07]$. We then evaluate three different objective functions (($L^{2}$), (KL), and (EE)), in analogous to the two-stream example. The results are displayed in Figure 5. Notably, in this one-dimensional setting, both ($L^{2}$) and (KL) exhibit non-convex behavior in the vicinity of the minimum, whereas (EE) displays a distinctly convex profile, mirroring the behavior observed in the two-stream example presented in Section 4.1.
Figure 5: Landscape of the Bump-on-Tail instability on the domain $[-0.07,0.07]$ for $b_{1}$ for (4.1). The objective functions are ($L^{2}$) (left), (KL) (center), and (EE)(right).
For Choice B, we extend the analysis to a two-parameter sweep by varying the pair $(b_{1},b_{2})$ over three pre-defined domains ($[-1,1]^{2}$, $[-0.1,0.1]^{2}$, and $[-0.001,0.003]^{2}$). The near-field region is shifted due to the new location of the global minimum in this case.
Figure 6 displays the values of the three objective functions (($L^{2}$), (KL), and (EE)) across the three domains. The conclusions drawn from the two-stream example continue hold, and in some cases even more pronounced. Notably, in the far-field region, the electric energy objective (EE), maintains a relatively convex structure and does not exhibit global minima for large values of $b_{1}$ and $b_{2}$. This observation indicates that an excessively strong external electric field does not always produce a correspondingly small self-generated electric field.
Figure 6: Landscape of the Bump-on-Tail instability with on the domain $[-1,1]^{2}$ (top), $[-0.1,0.1]^{2}$ (center) and, $[-0.001,0.003]^{2}$ (bottom) for $(b_{1},b_{2})$ for (4.2). The objective functions are ($L^{2}$) (left), (KL) (center), and (EE)(right).
@@ 5 Solving the PDE-constrained optimization problem
In this section, we present the numerical results obtained from solving (2.6) using Algorithm 1 to address the forward system described by (2.1). The computation was performed on an NVIDIA GeForce RTX 4090 GPU with 24 GB of memory. For the forward simulation, the phase space was discretized with $M=256$ mesh points, and the time domain was discretized using a time step of $\Delta t=0.1$. Each PDE solve can be completed within 5 seconds. Optimization was carried out using automatic differentiation, facilitated by the Python library JAX [jax2018github], to compute the required gradients.
The plots shown in Section 4 suggest that the landscape for the objectives ($L^{2}$) or (KL) are rather similar. Therefore we expect the performance of optimization solvers with these objectives would be similar as well. The rest of this section, we perform numerical studies only on (KL) and (EE).
The landscape plots only provide some suggestions on the performance of optimization solvers. Indeed, for local solvers such as gradient descent, that can only sees local behavior of objective functions, the solutions are trapped in local minimum, and thus placing the initial guess in the global basin is crucial. On the other hand, solvers have their non-convex extensions. One such possibility is to integrate the linesearch technique, which allows iterations to jump out of local minima. This can potentially relax the requirement for a good initial guess.
Another strategy typically deployed in optimization is “over-parameterization.” This is to increase the number of unknowns and adjustable parameters, and thus formulating the problem in a higher-dimensional space. In this setting, it is very likely that the dimension of the global optima manifold increases, and so the chance of obtaining optimal points.
@@
    • Local solver. There are many kinds of local solvers. We adopt the simplest gradient descent with constant stepsize. The stepsize is small enough ($81e-$ 8, $91e-$ 9) that one can essentially view the optimization iteration as a gradient flow along the landscape.
    • Adaptive/global solver. There are also many choices of adaptive solvers. To be more comparable with the choice of local solver (GD), we select GD with linesearch correction as our adaptive solver. The linesearch will satisfy the strong Wolfe condition that we briefly describe below.
@@
      Linesearch is imposed to ensure that the choice of $\alpha_{n}$ provides sufficient decrease. The “sufficiency” is quantified by the following:
@@
      for $10<c_{1}<c_{2}<$ 1. Namely, either the value of the new update is sufficiently smaller than the current value, or the new update has a gradient almost parallel to that of the old one. In particular, we use $c_{1}=10^{-4}$ and $c_{2}=0.9$.
  • Initial guess can be close, or far away from the optimum points;
    More specifically, the three kinds of initializations are
@@
      $H(x;b_{1},b_{2})=b_{1}\sin(k_{0}x)+b_{2}\sin(2k_{0}x)\,,$ (5.1)
@@
      $H(x;\bm{a},\bm{b})=\sum_{k=1}^{14}a_{k}\cos\left(kk_{0}x\right)+b_{k}\sin\left(kk_{0}x\right)$ (5.2)
@@ 5.1 Two Stream example
We summarize results for controlling two-stream instability, using the equilibrium state and initial condition presented in (2.9)-(2.10). All numerical results are presented in Table 1 that shows the final distribution function $f_{T}[H]$ with $H$ found by the optimization solver. Also, a more detailed summary of each experiment can be found in Appendix A.1. The stepsize used for the constant stepsize gradient descent is $10^{-8}$.
Under-parametrized Over-parametrized
Init. Step (KL) (EE) (KL) (EE)
Far Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(see Figure 7) (see Figure 13) (see Figure 19) (see Figure 25)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(see Figure 8) (see Figure 14) (see Figure 20) (see Figure 26)
Mid Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(see Figure 9) (see Figure 15) (see Figure 21) (see Figure 27)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(see Figure 10) (see Figure 16) (see Figure 22) (see Figure 28)
Near Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(see Figure 11) (see Figure 17) (see Figure 23) (see Figure 29)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(see Figure 12) (see Figure 18) (see Figure 24) (see Figure 30)
@@
  • Local solver with near-optimal initialization always stabilize the system, and such stability is achieved in both $2$-parameter regime, and $14$-parameter regime, as seen in Figures 121824 and 30. This signifies the importance of making a good initial guess. Making a good initial guess is usually not a feasible task for a classical PDE-constrained optimization. Fortunately, for plasma control, the study of the dispersion relation can potentially identify the global basin, see discussion in Section 3.
  • Adaptive solver is not compatible with the objective function (EE), as seen in Figures 13291527 and 29. Indeed, as shown in Figure 4, (EE) has very low evaluation when parameters are very large. When adaptive solvers such as GD with linesearch are deployed, the searching can take the iteration very far into the large parameter domain. This corresponds to the situation when the external field is extremely strong and it completely dominates the whole plasma behavior. When it happens, the self-generated electric field is indeed very small, but the solution renders meaningless physics. The only exception in this case is to set the initial data from a near-optimal initialization. The basin is convex enough to control the linesearch’s performance, see Figure 17.
  • The KL divergence does not seem to be a good metric for suppressing instability. Good suppression solution is only observed when initialization is near-optimal already (see Figures 1112 and 24). With neither adaptive or local solver, in neither under-parameterized or over-parameterized regime, KL metric does not return a satisfactory solution as long as initialization is not near-optimal.
  • Counter-intuitively, over-parametrization in this setting does not seem to bring too much benefit (comparing the third/forth columns with the first/second columns), with the only exception being the use of an adaptive solver for minimizing KL with a far-away initialization (comparing Figure 19 with Figure 7).
@@ 5.2 Bump-on-Tail example
For the Bump-on-Tail example (see Table 2 and Appendix A.2), most of the observations drawn from the previous example still hold true, or are even more prominent, except for the comparison between under- and over-parametrized scenarios. In particular, comparing over-parametrized results with those for under-parameterized using (EE) in the near-optimal initialization with local solver, it is evident over-parametrization significantly improves the stability (see Figures 54 and 42). This can potentially be explained by the fact that Bump-on-Tail distribution has more fine features and thus requires more Fourier modes for a successful suppression of instabilities.
Under-parametrized Over-parametrized
Init. Step (KL) (EE) (KL) (EE)
Far Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(see Figure 31) (see Figure 37) (see Figure 43) (see Figure 49)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(see Figure 32) (see Figure 38) (see Figure 44) (see Figure 50)
Mid Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(see Figure 33) (see Figure 39) (see Figure 45) (see Figure 51)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(see Figure 34) (see Figure 40) (see Figure 46) (see Figure 52)
Near Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(see Figure 35) (see Figure 41) (see Figure 47) (see Figure 53)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(see Figure 36) (see Figure 42) (see Figure 48) (see Figure 54)

### v3

@@ 4.1 Two Stream example
For Choice A, with one parameter, we conduct a sweeping of PDE solutions by setting $b_{1}$ to be in the pre-defined range of [0.1,0.1][-0.1,0.1], and compute the evaluation of the four objective functions ((KL), (EE), (KLT) and (EET)). The result is plotted in Figure 5. Evidently, in this 1D situation, both (EE) and (KL) that only see the last frame of the solution give non-convex behavior in the neighborhood of the minimum, whereas (EET) and (KLT) that see the whole solution in time demonstrates a convex profile. This strongly suggests the information in-time is crucial in determining the convexity of the landscape of the objective functions.
Figure 5: Landscape of the Two Stream instability on the domain [0.07,0.07][-0.07,0.07] for $b_{1}$ for (4.1). The objective functions are (KL)(left), (EE)(center-left), (KLT)(center-right) and (EET)(right).
@@
  • Far field: $1,1]\times[-1,1]$;
  • Mid-range: $0.1,0.1]\times[-0.1,0.1]$;
  • Near field: $0.003,0.001]\times[-0.003,0.001]$.
The wordings (“Far field, Mid-range, Near field”) are selected based on the conclusions drawn from Section 3 of dispersion relation, where it was suggested the optimal $H$ should be in the range of $O(\varepsilon)$. In this case, considering (2.10), $\varepsilon=0.001$. In Figure 6, we plot the values for the four different objective functions in these three domains, and some conclusions are drawn:
  • The landscape close to the global minimum (near-field) is convex for all four objective functions.
  • In the mid-range set, both plots that only use last-frame information ((KL)-(EE)) present a pronounced bulge around the minimum that highlights a clear non-convex behavior along the first mode direction ($b_{1}$ variable), while the plots that integrate in time ((KLT)-(EET)) eliminate these bulge and present a convex objective landscape.
  • In the far field, both (KL) and (EE) are highly oscillatory, making the global minima hard to find. In comparison, the objective functions that integrates in time are much smoother. Comparing KL and electric energy, we also see that, in both final-frame and integrated-in-time case, objective functions defined by electric energy present minima for large values of $b_{2}$, suggesting large external electric field can also suppress instability. However, we should point out that this is a non-physical setting: The imposed electric field $H$ is large, and completely dominates the dynamics.
Figure 6: Landscape of the Two Stream instability with on the domain $1,1]^{2}$(top), $0.1,0.1]^{2}$(center) and, $0.003,0.001]^{2}$(bottom) for $(b_{1},b_{2})$ for (4.2). The objective functions are (KL)(left), (EE)(center-left), (KLT)(center-right) and (EET)(right).
@@ 4.2 Bump-on-Tail example
In this section, we present results for the Bump-on-Tail instability problem with the following two choices of parameter settings:
  • Choice C:
    $H(x)=a_{1}\cos(k_{0}x)\,$, (4.3)
  • Choice D:
    $H(x)=a_{1}\cos(k_{0}x)+b_{1}\sin(k_{0}x)\,.$ (4.4)
For Choice C, we perform a parameter sweep of the PDE solutions by varying $a_{1}$ over the pre-defined interval [0.1,0.1][-0.1,0.1]. We then evaluate four different objective functions ((KL), (EE), (KLT) and (EET)), in analogous to the Two Stream example. The results are displayed in Figure 7. Notably, in this one-dimensional setting, both (KL) and (EE) exhibit non-convex landscape, whereas (KLT) and (EET) display a convex profile, mirroring the behavior observed in the Two Stream example in Figure 5. This observation resonates with the situation for the Two Stream instability, and suggests information in-time is crucial in convexifying the objective functions.
Figure 7: Landscape of the Bump-on-Tail instability on the domain [0.07,0.07][-0.07,0.07] for $a_{1}$ for (4.3). The objective functions are (KL)(left), (EE)(center-left), (KLT)(center-right) and (EET)(right).
For Choice D, we extend the analysis to a two-parameter sweep by varying the pair $(a_{1},b_{1})$ over three pre-defined domains ($1,1]^{2}$, $0.1,0.1]^{2}$, and $0.001,0.003]^{2}$). The near-field region is shifted due to the new location of the global minimum in this case.
Figure 8 displays the values of the four objective functions ((KL), (EE), (EET) and (KLT)) across the three domains. The conclusions drawn from the Two Stream example continue to hold in the near-field and mid-range cases. In the far-field region, (EET) and (KLT) continue to be more convex than those of (KL) and (EE). Moreover, they do not exhibit global minima for large values of $a_{1}$ and $b_{1}$ as seen in Figure 6 for far-field region. This observation indicates that an excessively strong external electric field does not always produce a correspondingly small self-generated electric field.
Figure 8: Landscape of the Bump-on-Tail instability with on the domain $1,1]^{2}$(top), $0.1,0.1]^{2}$(center) and, $0.001,0.003]^{2}$(bottom) for $(a_{1},b_{1})$ for (4.4). The objective functions are (KL)(left), (EE)(center-left), (KLT)(center-right) and (EET)(right).
@@ 5 Solving the PDE-constrained optimization problem
In this section, we present the numerical results obtained from solving (2.6) using Algorithm 1 to address the forward system described by (2.1). The computation was performed on an NVIDIA GeForce RTX 4090 GPU with 24 GB of memory. For the forward simulation, the phase space was discretized with $M=256$ mesh points, and the time domain was discretized using a time step of $\Delta t=0.1$. Each PDE solve can be completed within 5 seconds. Optimization was carried out using automatic differentiation, facilitated by the Python library JAX [9], to compute the required gradients.
The plots shown in Section 4 suggest that the landscape for the objectives (KLT) or (EET) are rather similar. Therefore we expect the performance of optimization solvers with these objectives would be similar as well. The rest of this section, we perform numerical studies only on (KL), (EE) and (EET).
The landscape plots only provide some suggestions on the performance of optimization solvers. Indeed, for local solvers such as gradient descent, that can only see local behavior of objective functions, the solutions are trapped in local minima, and thus placing the initial guess in the global basin is crucial. On the other hand, solvers have their non-convex extensions. One such possibility is to integrate the line-search technique, which allows iterations to jump out of local minima. This can potentially relax the requirement for a good initial guess.
Another strategy typically deployed in optimization is “over-parameterization”. This is to increase the number of unknowns and adjustable parameters, and thus formulating the problem in a higher-dimensional space. In this setting, it is very likely that the dimension of the global optima manifold increases, and so the chance of obtaining optimal points.
@@
    • Local solver. There are many kinds of local solvers. We adopt the simplest gradient descent with constant stepsize. The stepsize is small enough (1e81e-8, 1e91e-9) that one can essentially view the optimization iteration as a gradient flow along the landscape.
    • Adaptive/global solver. There are also many choices of adaptive solvers. To be more comparable with the choice of local solver (GD), we select GD with line-search correction as our adaptive solver. The line-search will satisfy the strong Wolfe condition that we briefly describe below.
@@
      Line-search is imposed to ensure that the choice of $\alpha_{n}$ provides sufficient decrease. The “sufficiency” is quantified by the following:
@@
      for $0<c_{1}<c_{2}<1$. Namely, either the value of the new update is sufficiently smaller than the current value, or the new update has a gradient almost parallel to that of the old one. In particular, we use $c_{1}=10^{-4}$ and $c_{2}=0.9$.
  • Initial guess can be close, or far away from the optimum points. More specifically, the three kinds of initializations are
@@
      $H(x;b_{1},b_{2})=b_{1}\sin(k_{0}x)+b_{2}\sin(2k_{0}x)\,,\quad\text{(Two Stream)}$ (5.1)
      $H(x;a_{1},b_{1})=a_{1}\cos(k_{0}x)+b_{1}\sin(k_{0}x)\,,\quad\text{(Bump-on-Tail)}$ (5.2)
@@
      $H(x;\bm{a},\bm{b})=\sum_{k=1}^{14}\big{(}a_{k}\cos\left(kk_{0}x\right)+b_{k}\sin\left(kk_{0}x\right)\big{)}$ (5.3)
@@ 5.1 Two Stream example
We summarize results for controlling Two Stream instability, using the equilibrium state and initial condition presented in (2.9)-(2.10). All numerical results are presented in Table 1 that shows the final distribution function $f_{T}[H]$ with $H$ found by the optimization solver. A more detailed summary of each experiment can be found in Appendix B.1. The stepsize used for the constant stepsize gradient descent is $10^{-8}$.
Under-parametrized (5.1) Over-parametrized (5.3)
Init. Step (KL) ($E$) (EET) (KL) ($E$) (EET)
Far Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 10) (Fig. 16) (Fig. 22) (Fig. 28) (Fig. 34) (Fig. 40)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 11) (Fig. 17) (Fig. 23) (Fig. 29) (Fig. 35) (Fig. 41)
Mid Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 12) (Fig. 18) (Fig. 24) (Fig. 30) (Fig. 36) (Fig. 42)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 13) (Fig. 19) (Fig. 25) (Fig. 31) (Fig. 37) (Fig. 43)
Near Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 14) (Fig. 20) (Fig. 26) (Fig. 32) (Fig. 38) (Fig. 44)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 15) (Fig. 21) (Fig. 27) (Fig. 33) (Fig. 39) (Fig. 45)
@@
  • Local solver with near-optimal initialization (good initial guess) perform universally well, and the stability is achieved in both under-parameter regime, and over-parameter regime, as seen in Figures 15, 21, 27, 33, 39 and 45. This signifies the importance of making a good initial guess. Making a good initial guess is usually not a feasible task for a classical PDE-constrained optimization, but is possible in this context by deploying the study of the dispersion relation, see discussion in Section 3.
  • Adaptive solver is not compatible with the objective functions (EE) and (EET), as seen in Figures 22, 40, 18, 24, 36, 42, 38 and 44. This is already suggested by Figure 6, (EET) where non-physical local minima can be found when parameters take on very large values. When adaptive solvers such as GD with line-search are deployed, the searching can take the iteration very far into these large parameter regions. The found external field is extremely strong and it completely dominates the whole plasma behavior. When it happens, the self-generated electric field is indeed very small, but the solution renders meaningless physics. The only exception in this case is to set the initial data from a near-optimal initialization. The basin is convex enough to control the line-search’s performance, see Figures 26 and 20.
  • The KL divergence does not seem to be a good metric for suppressing instability. The solution with good suppression is only observed when initialization is near-optimal already (see Figures 1415 and 33). With neither adaptive or local solver, in neither under-parameterized or over-parameterized regime, KL metric return a satisfactory solution as long as initialization is not near-optimal.
  • Counter-intuitively, over-parametrization in this setting does not seem to bring too much benefit (comparing the last three columns with the first three columns), with the only exception being the use of an adaptive solver for minimizing KL with a far-away initialization (comparing Figure 28 with Figure 10).
  • Without good initial guess, the only good solution is found by optimizing (EET) with simple GD solver in the properly parametrized setting when initialization is prepared in the mid-range (see Figure 25).
@@ 5.2 Bump-on-Tail example
For the Bump-on-Tail example, we conduct the same experiments and summarize them in Table 2 and Appendix B.2. Most observations drawn from the previous example still hold true. Yet we should identify the distinct features. In particular, our simulation seems to suggest that Bump-on-Tail instability is easier to be controlled, though the perfect recovery is more rare. More specifically, for Two Stream instability, when we have good initial guess, we obtained exact recovery when we use adaptive solvers in the under-parameterized regime and local solvers in the over-parameterized regime. We fail to do so in the Bump-on-Tail situation except for (EET). On the other hand, we find many more situations in which, though we are not being able to find exact recovery, the solutions are still relatively stable. This strongly suggests the local minima are good controls nevertheless, see trajectory data in Figures 46, 47, 59, 53, 51, 63 and 57.
Finding good initialization is still instrumental and it brings out-performance across solvers and regimes. Adaptive solvers are still incompatible with (EE) and (EET), and the performance is universally bad. Over-parameterization still does not necessarily bring benefits. In particular, with good initial guess and local solver, looking for best control in the over-parameterized regime brings even worse reconstruction, see Figures 69 and 81 vs. Figures 51 and 63.
However, in comparison to the Two Stream case, KL divergence seems to be a reasonable objective function, by minimizing which, acceptable controls for Bump-on-Tail can still be found. This resonates our earlier discussion on local minima serving as good controls for Bump-on-Tail instability.
Under-parametrized (5.2) Over-parametrized (5.3)
Init. Step (KL) ($E$) (EET) (KL) ($E$) (EET)
Far Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 46) (Fig. 52) (Fig. 58) (Fig. 64) (Fig. 70) (Fig. 76)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 47) (Fig. 53) (Fig. 59) (Fig. 65) (Fig. 71) (Fig. 77)
Mid Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 48) (Fig. 54) (Fig. 60) (Fig. 66) (Fig. 72) (Fig. 78)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 49) (Fig. 55) (Fig. 61) (Fig. 67) (Fig. 73) (Fig. 79)
Near Adaptive [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 50) (Fig. 56) (Fig. 62) (Fig. 68) (Fig. 74) (Fig. 80)
Local [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(Fig. 51) (Fig. 57) (Fig. 63) (Fig. 69) (Fig. 75) (Fig. 81)